OpenStudy (anonymous):

Find lim (sin(root of (x+h))-sin(root of (x)) / h as h goes to 0 please I need help to solve this problem

hartnn (hartnn):

and you can't use L'Hopital's rule?

hartnn (hartnn):

If not, then using the sin A - sin B formula could be a good start, try it :)

OpenStudy (anonymous):

yes, i agree

OpenStudy (anonymous):

http://www.purplemath.com/modules/idents.htm sum identities

OpenStudy (pawanyadav):

I think x is a constant here so you can use L'hopital rule So diff. of sin√(x+h)=cos√(x+h)/2√(x+h) And diff. Of Sin√x =0 as x an constant So you get 1 in denominator also Then answers will be Cos√(x+h)/2√(x+h) You have now just to put h=0 Then what Will you get ??

OpenStudy (anonymous):

Thank youuuuuuu so much

OpenStudy (pawanyadav):

You are welcome